# Weak interaction axial form factors of the octet baryons in nuclear medium

### v1

@@ IV Extension of the model to symmetric nuclear medium
Figure 3: Axial-vector form factor $G_{A}$ for $\Delta I=1$ transitions. Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
The axial-vector form factors of baryon octet in vacuum was already calculated in Ref. [8]. We consider now the extension to the nuclear medium. An important property observed in the nuclear medium is the suppression of the axial-vector form factors due to the enhancement of the $P$-state component, in general associated with the lower component of the quark Dirac spinors. In the present case the simplest way to include the $P$-state effect is to modify the value of $n_{P}$. In the absence of a solid, established reference, we adjust the value of $n_{P}$ in order to reproduce the suppression estimated by the Bag Model from Ref. [78]. Since our bare estimate for $G_{A}^{\rm B}(0)$ for the nucleon is similar to the Bag Model $g_{A}^{N}$, we use the in-medium to free space bare axial-vector coupling ratio $g_{A}^{N\ast}/g_{A}^{N}$ to fix the amount of $P$-state admixture (coefficient $n_{P}$). Notice that the determination of $n_{P}$ by the values of $g_{A}^{N\ast}/g_{A}^{N}$ is not an additional constraint of the model. It is a necessary condition associated with the consistency with the QMC model. The QMC model uses the input of the Bag Model for $g_{A}^{N\ast}$ [78] and this result is used in the calculation of coupling constants $g^{\ast}_{MB_{1}B_{2}}$ [28]. When we include the effect of the meson cloud for the nucleon on the functions $G_{A}^{\ast}$ we modify the values of $G_{A}(0)$ and $G_{A}^{\ast}(0)$ and consequently the ratio $G_{A}^{\ast}(0)/G_{A}(0)$ deviates from the $g_{A}^{N\ast}/g_{A}^{N}$.
The determination of the value of $n_{P}$ is based on the relation [8]
$\tilde{G}_{A}^{\rm B}(0)=\frac{15-19n_{P}^{2}}{9}g_{A}^{q}(0),$ (37)
where $g_{A}^{q}(0)=1$.
@@
Figure 4: Axial-vector form factor $G_{A}$ for $\Delta S=1$ transitions (part 1). Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
$\rho/\rho_{0}$ $M_{N}$ $M_{\Lambda}$ $M_{\Sigma}$ $M_{\Xi}$ $m_{K}$ $m_{\rho}$ $m_{\phi}$ $g_{\pi NN}$ $n_{P}$
0 0.9390 1.1160 1.1920 1.3180 0.4937 0.7545 1.0195 1.000 $-0.507$
0.5 0.8313 1.0439 1.1214 1.2822 0.4573 0.7061 1.0191 0.921 $-0.536$
1.0 0.7545 0.9927 1.0704 1.2567 0.4305 0.6537 1.0189 0.899 $-0.560$
Table 6: Model parameters used in the extension to the nuclear medium (baryon octet case). The discussion about the values of $n_{P}$ are in the previous sections. The kaon mass is used in the calculation of the $G_{P}$ pole term in the $\Delta S=1$ transitions. For $m_{\pi}^{*}\simeq m_{\pi}$ we use $m_{\pi}=0.138$ GeV [17, 18, 19, 20].
V Numerical results
In this section we present our numerical results for the octet baryon axial form factors $G_{A}$ and $G_{P}$ in vacuum and in medium.
We start with the discussion of the results for the vacuum and some remarks about the calibration of our model in vacuum (Sec. V.1). Next we discuss the results in the nuclear medium (Sec. V.2). We display the vacuum and in-medium results side by side in order to better observe the modifications due to the nuclear medium (Figs. 3, 4 and 5 and Figs. 6, 7 and 8). We use also the ratio between the form factors in medium and vacuum for a clear visualization of the medium modifications. Deviations from the unity (one) are the signature of medium effects.
We study also the slope of the form factors near $Q^{2}=0$ (Sec. V.3). We finish the section with a general discussion of our results, the limitations of the model, and the possible improvements (Sec. V.4).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 5: Axial-vector form factor $G_{A}$ for $\Delta S=1$ transitions (part 2). Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 6: Induced pseudoscalar form factor $G_{P}$ for $\Delta I=1$ transitions. Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$).
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 7: Induced pseudoscalar form factor $G_{P}$ for $\Delta S=1$ transitions (part 1). Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). Same ratios for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 8: Induced pseudoscalar form factor $G_{P}$ for $\Delta S=1$ transitions (part 2). Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). Same ratios for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$.
@@ V.1 Axial form factors in vacuum
The results of the $B^{\prime}\to B$ octet baryon axial-vector form factors $G_{A}$ are presented on the each left side of Figs. 3, 4 and 5. The contributions from the quark core are represented by the dashed lines and the contributions from the meson cloud by the dashed-doted lines. The final results correspond to the solid lines. In the cases where the final result $G_{A}$ is negative, we represent the function $-G_{A}$, for an easier comparison with the other reactions. In the figures we include also the available data for $G_{A}(0)$.
From the figures, we can conclude that we can have different relative magnitudes for the form factor, for the bare contribution and for the meson cloud contribution. Most transitions are dominated by the bare contributions but some transitions have significant relative meson cloud contributions ($\Xi^{-}\to\Xi^{0}$, $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{0}\to\Sigma^{+}$). Transitions with a large magnitude ($|G_{A}(0)|\gtrsim 1$) can have about 33% of meson cloud contributions (transitions $n\to p$, $\Lambda\to p$, $\Xi-\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$).
In comparison with the previous work [8] we correct here the normalization of the $\Xi$ baryons, relatively to the previous work, as discussed already in Sec. III. The main difference to the previous calculations are then the results for $\Xi^{-}\to\Lambda$ and $\Xi^{0}\to\Sigma^{+}$, which are modified by the value of $Z_{\Xi}$. As a consequence, both, the bare and the meson cloud contributions are enhanced. With the present parametrization the description of $\Xi^{-}\to\Lambda$ transition is improved, while the numerical result for $\Xi^{0}\to\Sigma^{+}$ overestimates the data.
@@
Figure 9: Axial-vector form factor $G_{A}$ for $\Delta I=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 10: Axial-vector form factor $G_{A}$ for $\Delta S=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$. Same ratios for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$. Same ratios for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 11: Induced pseudoscalar form factor $G_{P}$ for $\Delta I=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$.
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Refer to caption           Refer to caption
Figure 12: Induced pseudoscalar form factor $G_{P}$ for $\Delta S=1$ transitions. Ratios $G_{P}^{\ast}/G_{P}$ and $G_{P}^{\ast}(Q^{2})/G_{P}^{\ast}(Q_{p}^{2})$. Same ratios for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$. Same ratios for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$.
@@
Our predictions for $G_{P}$ are on the each left side of Fig. 6 for $\Delta I=1$, and on the each left side of Figs. 7, 8 for $\Delta S=1$. In the figures, we discriminate the result for the bare contribution, the combination of the bare and pseudoscalar meson pole contribution, and the sum of all the contributions (Bare + Pole + Meson Cloud). We adjust the upper limits of $Q^{2}$ in order to discriminate the different contributions at low $Q^{2}$ according to the falloff of the form factors. As for the function $G_{A}$, in the cases where the final result $G_{P}$ is negative, we represent the function $-G_{P}$.
In the calculations for $\Delta I=1$ (Fig. 6) it is clear the impact of the factor $1/(m_{\pi}^{2}+Q^{2})$ in the Bare + Pole result and in the final result. The strong reduction with $Q^{2}$ is a consequence of the small magnitude of $m_{\pi}$. In the case of the nucleon, the combination of the bare contribution and the meson cloud contribution is decisive for the good agreement of the model with the physical form factor data (see Fig. 2 and Sec. III). In the remaining cases our calculations are pure predictions.
The results for $\Delta S=1$ (Figs. 7 and 8) show a slower falloff for the form factor $G_{P}$ than for $\Delta I=1$, as expected from the kaon pole contribution associated with the factor $1/(m_{K}^{2}+Q^{2})$. The magnitude of $G_{P}$ decreases slower with $Q^{2}$ than for $\Delta I=1$ as a consequence of the kaon mass of the pole. The differences of magnitudes near $Q^{2}=0$ in comparison with $\Delta I=1$ are mainly the consequence of the amplification of the meson mass.
Another important point is the relative magnitude of the terms. The bare and pole terms have, in the case $\Delta S=1$, similar magnitudes, and there is a partial cancellation between the pole terms and the bare terms, which have opposite signs for all transitions. Notice that, since the pole and meson cloud contributions have the same sign for the transitions under discussion, one can also say that there is a partial cancellation between bare and meson cloud contributions at low $Q^{2}$. The consequence of these effects is that the form factors $G_{P}$ for $\Delta S=1$ are about one order of magnitude smaller than the form factors $G_{P}$ for $\Delta I=1$. In a few cases ($\Xi^{-}\to\Lambda$, $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$) the bare contribution has a magnitude comparable with the meson cloud contribution at low $Q^{2}$.
@@
Our last remark about $G_{P}$ is the enhancement of $G_{P}$ near $Q^{2}=0$ as a consequence of the low-$Q^{2}$ turning point of the function $G_{P}^{\rm B}$. The effect appears also for $\Delta I=1$, but it is overshadowed by the magnitude of the pole and meson cloud terms. It can be observed more clearly in the figures for $\Delta S=1$. This behavior is a consequence of the method used to calibrate the model: an unconstrained fit to the lattice QCD data for the nucleon. The lattice data sets used in the calibration include no $G_{P}$ data below $Q^{2}=0.15$ GeV2 [38]. Therefore, it is possible that $G_{P}^{\rm B}$ may be poorly constrained, affecting consequently the accuracy of the description of $G_{P}$ near $Q^{2}=0$. We will back to this issue in the final discussion (Sec. V.4).
In the present study, we use the calibration from Ref. [8] as a first approximation. The results are in principle more accurate for the $\Delta I=1$ sector, less sensitive to the magnitude of the function $G_{P}^{\rm B}$. Future improvements of the present scheme are discussed in Sec. V.4.
@@ V.2 Axial form factors in medium
The result of our extension of the model to the octet baryons in the nuclear medium are presented on the each right side of Figs. 3, 4 and 5 for $G_{A}$ and on the each right side of Figs. 6, 7 and 8 for $G_{P}$. In the calculations we use the parameters from Table 6.
For the nuclear medium, for simplicity, we present only the final results for the densities $\rho=0.5\rho_{0}$, $\rho=\rho_{0}$, in comparisons with the result in vacuum ($\rho=0$), omitting the decomposition of the bare and meson cloud contributions.
In the discussion of the results in medium we also discuss the ratios $G_{A}^{\ast}/G_{A}$ presented on the each left side of Figs. 9 and 10 and the ratios $G_{P}^{\ast}/G_{P}$ presented on the each left side of Figs. 11 and 12. In this respect, we recall the discussion of Sec. III.3, according to, the ratios for $\Sigma^{0}\to p$ are the same as the ratios for $\Sigma^{-}\to n$. The same property is verified for the transitions $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$.
@@ V.2.1 Results for $G_{A}^{\ast}$
From the observation of the each right side of Figs. 3, 4 and 5, and the each left side of Figs. 9 and 10, we conclude that, in general, the axial-vector form factor $G_{A}$ is reduced in the nuclear medium, or symmetric nuclear matter, (suppression of $G_{A}$ in medium). The conclusion is valid for $\Delta I=1$ and $\Delta S=1$.
@@
In the present case these results are as a consequence of the suppression in medium of both the meson cloud and bare contributions. The suppression of the meson cloud contribution is more effective at low $Q^{2}$, while the suppression of the bare contributions is more effective at large $Q^{2}$. The magnitude of the suppression can be better observed using the function $G_{A}^{\ast}/G_{A}$ in terms of $Q^{2}$, presented on the left side of Fig. 9 for $\Delta I=1$ and on the left side of Fig. 10 for $\Delta S=1$. The suppression is stronger for lighter baryons and larger densities.
From the observation of Figs. 9 and 10, we can classify the ratios $G_{A}^{\ast}/G_{A}$ using the shape and the magnitude of the suppression in the order:
@@
The range of variation discussed next are relative to the interval $Q^{2}=0$–2 GeV2.
The first reaction ($n\to p$), includes no transitions between the strange and light quarks, and thus it is the reaction with stronger suppression in medium. The effect increases with $Q^{2}$. For $\rho/\rho_{0}=0.5$ one has a reduction of about 5–18%. For $\rho/\rho_{0}=1.0$ the reduction varies from 8–32%.
The second type includes the $\Lambda\to p$ transition, that is similar to the $n\to p$ transition, but with a weaker reduction of the ratio $G_{A}^{\ast}/G_{A}$. The $\Sigma^{+}\to p$/$\Sigma^{-}\to n$ ratio have a similar shape.
The third class of transitions ($\Sigma^{+}\to\Lambda$, $\Sigma\to\Sigma,\Lambda$, $\Xi\to\Sigma,\Lambda$) includes systems composed by a strange quark in the final state and have smaller and slower reductions with $Q^{2}$.
The last case ($\Xi^{-}\to\Xi^{0}$) is a transition between baryons with two strange quarks and reveal a significant suppression at low $Q^{2}$ (meson cloud contributions significantly reduced in medium) and only a mild dependence on $Q^{2}$, for $Q^{2}>1$ GeV2.
Our estimates for $G_{A}^{\ast}/G_{A}$ for the nucleon can be compared with the calculations of the Bag Model [78, 25, 26]. The main difference between our calculations and the Bag Model calculations is that the latter does not include the effects of the meson cloud dressing. Apart small differences at low $Q^{2}$, the calculations differ on the shape of the ratio. The Bag Model ratio estimates decrease till the region $Q^{2}=0.5$–1 GeV2, and start to increase after that. After a certain value of $Q^{2}$, one has an enhancement of $G_{A}^{\ast}$. This is due to the Lorentz contraction effect that is taken into account for the ”bag shape” in Ref. [78], and thus for larger $Q^{2}$ the effect starts to appear stronger. Or, this is due to the finite size effect of the nucleon bag which is subject to the Lorentz contraction. The present calculations in contrast, suggests that the medium suppression keeps increasing for large $Q^{2}$ although with a reduced rate.

### v2

@@ V Numerical results
Figure 4: Axial-vector form factor $G_{A}$ for $|\Delta S|=1$ transitions. Left panel: results in vacuum (bare, meson cloud and total). Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). The results for the $\Sigma^{-}\to n$ transition are presented in Appendix B, Fig. B2. The results for $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{0}$ transitions are presented in Appendix B, Fig. B3.
We use the ratio between the form factors in medium and vacuum for a clear visualization of the medium modifications (Sec. V.3). Deviations from unity (one) are the signature of medium effects. We study also the slope of the form factors near $Q^{2}=0$.
We summarize the results for $G_{A}$ and $G_{P}$ in medium in Sec. V.4. We finish the section with a general discussion of our results, the limitations of the model, and the possible improvements (Sec. V.5).
@@
Refer to caption           Refer to caption
Figure 5: Induced pseudoscalar form factor $G_{P}$ for $|\Delta I|=1$ transitions. Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). The results for the $\Sigma^{+}\to\Lambda$ and $\Sigma^{0}\to\Sigma^{+}$ transitions are presented in Appendix C, Fig. C1.
@@
Figure 6: Induced pseudoscalar form factor $G_{P}$ for $|\Delta S|=1$ transitions. Left panel: results for bare, bare plus meson cloud and total. Right panel: total results for the medium $\rho=0.5\rho_{0}$ and $\rho_{0}$ compared with vacuum ($\rho=0$). The results for the $\Sigma^{-}\to n$ transition are presented in Appendix C, Fig. C2. The results for $\Xi^{-}\to\Lambda$ and $\Xi^{-}\to\Sigma^{0}$ transitions are presented in Appendix C, Fig. C3.
@@ V.1 Axial form factors in vacuum
Some examples of the $B^{\prime}\to B$ octet baryon axial-vector form factors $G_{A}$ are presented on the left side in Figs. 3 ($|\Delta I|=1$) and 4 ($|\Delta S|=1$). The results for all transitions are presented in Appendix B (Figs. B1, B2 and B3). The contributions from the quark core are represented by the dash lines and the contributions from the meson cloud by the dash-doted lines. The final results correspond to the solid lines. In the cases where the final result $G_{A}$ is negative, we represent the function $-G_{A}$, for an easier comparison with the other reactions. In the figures we include also the available data for $G_{A}(0)$.
From these figures and from Figs. B1, B2 and B3, we can conclude that we can have different relative magnitudes for the form factor, for the bare contribution and for the meson cloud contribution. Most transitions are dominated by the bare contributions but some transitions have significant relative meson cloud contributions ($\Xi^{-}\to\Xi^{0}$, $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{0}\to\Sigma^{+}$). Transitions with a large magnitude ($|G_{A}(0)|\gtrsim 1$) can have about 33% of meson cloud contributions (transitions $n\to p$, $\Lambda\to p$, $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$).
In comparison with previous work [7] we correct here the normalization of the $\Xi$ baryons, relatively to the previous work, as discussed already in Sec. III. The main difference to the previous calculations are then the results for $\Xi^{-}\to\Lambda$ and $\Xi^{0}\to\Sigma^{+}$, which are modified by the value of $Z_{\Xi}$. As a consequence, both, the bare and the meson cloud contributions are enhanced. With the present parametrization the description of $\Xi^{-}\to\Lambda$ transition is improved, while the numerical result for $\Xi^{0}\to\Sigma^{+}$ overestimates the data.
@@
Figure 7: Axial-vector form factor $G_{A}$ for $|\Delta I|=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$. The results for the $\Sigma^{+}\to\Lambda$ and $\Sigma^{0}\to\Sigma^{+}$ transitions are presented in Appendix B, Fig. B4.
@@
Figure 8: Axial-vector form factor $G_{A}$ for $|\Delta S|=1$ transitions. Ratios $G_{A}^{\ast}/G_{A}$ and $G_{A}^{\ast}(Q^{2})/G_{A}^{\ast}(0)$. The results for the $\Sigma^{-}\to n$, $\Sigma^{0}\to p$ and $\Xi^{-}\to\Lambda$ transition are presented in Appendix B, Fig. B5. The ratios are the same for $\Sigma^{-}\to n$ and $\Sigma^{0}\to p$. The ratios are the same for $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$.
@@
Some examples of our predictions for $G_{P}$ are on the left side of Fig. 5 for $|\Delta I|=1$, and on the left side of Fig. 6 for $|\Delta S|=1$. The results for $G_{P}$ for all transitions are presented in Appendix C (Figs. C1, C2 and C3). As for the function $G_{A}$, in the cases where the final result $G_{P}$ is negative, we represent the function $-G_{P}$. In the figures, we discriminate the result for the bare contribution, the combination of the bare and pseudoscalar meson pole contribution, and the sum of all the contributions (Bare + Pole + Meson Cloud). We adjust the upper limits of $Q^{2}$ in order to discriminate the different contributions at low $Q^{2}$ according to the falloff of the form factors.
In the calculations for $|\Delta I|=1$ (Figs. 5 and C1) it is clear the impact of the factor $1/(m_{\pi}^{2}+Q^{2})$ in the Bare + Pole result and in the final result. The strong reduction with $Q^{2}$ is a consequence of the small magnitude of $m_{\pi}$. In the case of the nucleon, the combination of the bare contribution and the meson cloud contribution is decisive for the good agreement of the model with the physical form factor data (see Fig. 2 and Sec. III). In the remaining cases our calculations are pure predictions.
The results for $|\Delta S|=1$ (Figs. 6, C2 and C3) show a slower falloff for the form factor $G_{P}$ than for $|\Delta I|=1$, as expected from the kaon pole contribution associated with the factor $1/(m_{K}^{2}+Q^{2})$. The magnitude of $G_{P}$ decreases slower with $Q^{2}$ than for $|\Delta I|=1$ as a consequence of the kaon mass of the pole. The differences of magnitudes near $Q^{2}=0$ in comparison with $|\Delta I|=1$ are mainly the consequence of the amplification of the meson mass.
Another important point is the relative magnitude of the terms. The bare and pole terms have, in the case $|\Delta S|=1$, similar magnitudes, and there is a partial cancellation between the pole terms and the bare terms, which have opposite signs for all transitions. Notice that, since the pole and meson cloud contributions have the same sign for the transitions under discussion, one can also say that there is a partial cancellation between bare and meson cloud contributions at low $Q^{2}$. The consequence of these effects is that the form factors $G_{P}$ for $|\Delta S|=1$ are about one order of magnitude smaller than the form factors $G_{P}$ for $|\Delta I|=1$. In a few cases ($\Xi^{-}\to\Lambda$, $\Xi^{-}\to\Sigma^{0}$ and $\Xi^{0}\to\Sigma^{+}$) the bare contribution has a magnitude comparable with the meson cloud contribution at low $Q^{2}$.
@@
Our last remark about $G_{P}$ is the enhancement of $G_{P}$ near $Q^{2}=0$ as a consequence of the low-$Q^{2}$ turning point of the function $G_{P}^{\rm B}$. The effect appears also for $|\Delta I|=1$ (Figs. 5 and C1), but it is overshadowed by the magnitude of the pole and meson cloud terms. It can be observed more clearly in the figures for $|\Delta S|=1$ (Figs. 6, C2 and C3). This behavior is a consequence of the method used to calibrate the model: an unconstrained fit to the lattice QCD data for the nucleon. The lattice datasets used in the calibration include no $G_{P}$ data below $Q^{2}=0.15$ GeV2 [38]. Therefore, it is possible that $G_{P}^{\rm B}$ may be poorly constrained, affecting consequently the accuracy of the description of $G_{P}$ near $Q^{2}=0$.
In the present study, we use the calibration from Ref. [7] as a first approximation. The results are, in principle, more accurate for the $|\Delta I|=1$ sector, less sensitive to the magnitude of the function $G_{P}^{\rm B}$. Future improvements of the present scheme are discussed in Sec. V.5.
@@ V.2 Axial form factors in medium
Some examples of our extension of the model to the octet baryons in the nuclear medium are presented on the right side of Figs. 3 ($|\Delta I|=1$) and 4 ($|\Delta S|=1$) for $G_{A}$, and on the right side of Figs. 5 ($|\Delta I|=1$) and 6 ($|\Delta S|=1$) for $G_{P}$. The results for all transitions are presented in Appendix B (Figs. B1, B2 and B3) for $G_{A}$, and Appendix C (Figs. C1, C2 and C3) for $G_{P}$.
For the nuclear medium, for simplicity, we present only the final results for the densities $\rho=0.5\rho_{0}$ and $\rho=\rho_{0}$, in comparisons with the result in vacuum ($\rho=0$), omitting the decomposition of the bare and meson cloud contributions.
For the discussion of the results in medium it is important to look to the ratios $G_{A}^{\ast}/G_{A}$ presented on the left side of Figs. 7 and 8 and the ratios $G_{P}^{\ast}/G_{P}$ presented on the left side of Figs. 9 and 10. The ratios for all transitions are presented in Appendix B (Figs. B4 and B5) for $G_{A}$, and in Appendix C (Figs. C4 and C5) for $G_{P}$. In this respect, we recall the discussion in Sec. III.3, according to which the ratios for $\Sigma^{0}\to p$ are the same as the ratios for $\Sigma^{-}\to n$. The same property is verified for the transitions $\Xi^{0}\to\Sigma^{+}$ and $\Xi^{-}\to\Sigma^{0}$.
@@ V.2.1 Results for $G_{A}^{\ast}$
From the results for the form factor $G_{A}$ in medium: right side of Figs. 3, 4, B1, B2 and B3 for the function $G_{A}^{\ast}$, and left side of Figs. 7, 8, B4 and B5 for the function $G_{A}^{\ast}/G_{A}$, we conclude that, in general, the axial-vector form factor $G_{A}$ is reduced in the nuclear medium, or symmetric nuclear matter (suppression of $G_{A}$ in medium). The conclusion is valid for $|\Delta I|=1$ and $|\Delta S|=1$.
@@
In the present case these results are a consequence of the suppression in medium of both the meson cloud and bare contributions. The suppression of the meson cloud contribution is more effective at low $Q^{2}$, while the suppression of the bare contributions is more effective at large $Q^{2}$. The magnitude of the suppression can be better observed using the function $G_{A}^{\ast}/G_{A}$ in terms of $Q^{2}$ (left side of Figs. 7 and B4 for $|\Delta I|=1$, and left side of Figs. 8 and B5 for $|\Delta S|=1$). The suppression is stronger for lighter baryons and larger densities.
From the observation of Figs. 7, 8, B4 and B5, we can classify the ratios $G_{A}^{\ast}/G_{A}$ using the shape and the magnitude of the suppression in the order:
@@
The range of variation discussed next is relative to the interval $Q^{2}=0$–2 GeV2.
The first reaction ($n\to p$) includes no transitions between the strange and light quarks, and, thus, it is the reaction with stronger suppression in medium. The effect increases with $Q^{2}$. For $\rho/\rho_{0}=0.5$ one has a reduction of about 5–18%. For $\rho/\rho_{0}=1.0$ the reduction varies from 8–32%.
The second type includes the $\Lambda\to p$ transition, that is similar to the $n\to p$ transition, but with a weaker reduction of the ratio $G_{A}^{\ast}/G_{A}$. The $\Sigma^{+}\to p$/$\Sigma^{-}\to n$ ratio has a similar shape.
The third class of transitions ($\Sigma\to\Lambda,\Sigma$, $\Xi\to\Lambda,\Sigma$) includes systems composed by a strange quark in the final state and have smaller and slower reductions with $Q^{2}$.
The last case ($\Xi^{-}\to\Xi^{0}$) is a transition between baryons with two strange quarks and reveals a significant suppression at low $Q^{2}$ (meson cloud contributions significantly reduced in medium) and only a mild dependence on $Q^{2}$, for $Q^{2}>1$ GeV2.
Our estimates for $G_{A}^{\ast}/G_{A}$ for the nucleon can be compared with the calculations of the bag model [66, 25, 26]. The main difference between our calculations and the bag model calculations is that the latter does not include the effects of the meson cloud dressing. Apart from small differences at low $Q^{2}$, the calculations differ on the shape of the ratio. The bag model ratio estimates decrease till the region $Q^{2}=0.5$–1 GeV2, and start to increase after that. After a certain value of $Q^{2}$, one has an enhancement of $G_{A}^{\ast}$. This is due to the Lorentz contraction effect that is taken into account for the ”bag shape” in Ref. [66], and, thus, for larger $Q^{2}$ the effect starts to appear stronger. Or, this is due to the finite size effect of the nucleon bag which is subject to the Lorentz contraction. The present calculations in contrast, suggest that the medium suppression keeps increasing for large $Q^{2}$ although with a reduced rate.